When the astronauts left the Moon, they used a spacecraft with a mass of 2150 kg 

Calculate the height reached by the spacecraft at the point where it had a  gravitational potential energy of 86 000 000 J 

The gravitational field strength of the Moon is 1.6 N/kg 


1
Expert's answer
2021-02-28T07:23:36-0500

By definition, the gravitational potential enegry is given as follows:


"W = mgh"

where "m = 2150kg" is the mass of the spacecraft, "g = 1.6N\/kg" is the gravitational field strength, and "h" is the height reached by the spacecraft. SInce "W = 86 000 000 J", obtain:


"h = \dfrac{W}{mg}\\
h = \dfrac{86 000 000 J}{2150kg\cdot 1.6N/kg} =25000m = 25km"

Answer. 25 m or 25 km.